# Make Your Guard Learn to Think: Defending Against Jailbreak Attacks with Safety Chain-of-Thought

### v1

@@ Abstract
Large language models (LLMs) are vital for a wide range of applications yet remain susceptible to jailbreak threats, which could lead to the generation of inappropriate responses. Conventional defenses, such as refusal and adversarial training, often fail to cover corner cases or rare domains, leaving LLMs still vulnerable to more sophisticated attacks. We propose a novel defense strategy, Safety Chain-of-Thought (SCoT), which harnesses the enhanced reasoning capabilities of LLMs for proactive assessment of harmful inputs, rather than simply blocking them. SCoT augments any refusal training datasets to critically analyze the intent behind each request before generating answers. By employing proactive reasoning, SCoT enhances the generalization of LLMs across varied harmful queries and scenarios not covered in the safety alignment corpus. Additionally, it generates detailed refusals specifying the rules violated. Comparative evaluations show that SCoT significantly surpasses existing defenses, reducing vulnerability to out-of-distribution issues and adversarial manipulations while maintaining strong general capabilities. The code and data is available at https://anonymous.4open.science/r/SCoT-D4D9.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) exhibit exceptional capabilities, enabling their wide applications in fields such as education (zhang2024simulating), programming (perez2020copilot), and everyday tasks. However, their powerful nature also introduces risks, as they can inadvertently generate harmful instructions or inappropriate content, leading to unsafe or illegal outcomes. For instance, prior research has shown that malicious users can compromise LLMs through adversarial techniques such as jailbreaks (zou2023universal; liu2024autodan), which bypass safety restrictions and prompt models to generate harmful outputs. These outputs may include instructions for breaching computer systems (zou2023universal) or facilitating unauthorized access to copyrighted materials (mazeika2024harmbenchstandardizedevaluationframework). Such misuse not only violates ethical standards in AI deployment but also leads to severe consequences, including financial losses, reputational harm, and legal liabilities. Given these concerns, ensuring the safety and ethical deployment of LLMs is paramount. Addressing safety challenges is not only critical to preventing potential harm but also essential to fostering trust and acceptance of these technologies in society.
@@
A common way to achieve safety defense is safety alignment or refusal training (bai2022traininghelpfulharmlessassistant), where LLMs are trained to reject harmful instructions. However, these strategies fail to defend against increasingly sophisticated attacks as their design and training recipes are still vulnerable to these emerging threats. These attacks often take advantage of two main weaknesses of the refusal training: (1) out-of-distribution scenarios, where the harmful input is not covered by safety training datasets but is in the pretraining stage, and (2) competing objectives, where attackers could add harmful suffixes or introduce distracting role-play content to harmful questions to manipulate the model into starting its response positively (wei2024jailbroken). These methods exploit the way LLMs are trained and predict text, leading to inappropriate outputs. Even advanced models like GPT-4 (openai2024gpt4technicalreport) and Claude 3.5 Sonnet (Anthropic2024) remain vulnerable to these sophisticated methods, underscoring the urgent need for more adaptive and robust defense mechanisms.
We envision that human’s thinking process of responding to a given harmful query is to access whether it is appropriate to give the answer in mind and then provide the response to the answer or refusal. Motivated by this cognitive process humans approach complex decisions and the recent advancements in the reasoning abilities of LLMs, we propose a novel defense strategy that fundamentally diverges from traditional approaches. This method, which we name SCoT: Safety Chain-of-Thought, instructs the model to first analyze the potential harmfulness of the request before giving the corresponding response. As depicted in Figure 1, this reasoning-based approach allows the model to engage in a “thinking process” to categorize the request into predefined violations, enabling it to generalize effectively to different variants of harmful questions and out-of-distribution scenarios that are not explicitly covered in the safety training corpus. Moreover, this proactive reflective mechanism naturally defends against attacks designed to elicit affirmative outputs. By requiring the model to evaluate the request’s intent before responding, our SCoT mitigates the risk of manipulation, ensuring a higher degree of robustness and adaptability compared to traditional defenses.
We rigorously evaluate SCoT across a comprehensive range of attack scenarios, encompassing 10 distinct black-box and white-box attack types. SCoT is benchmarked against safety-aligned models and some of the most advanced state-of-the-art defense strategies as outlined in recent studies (zou2024circuitbreaker; mazeika2024harmbenchstandardizedevaluationframework). Comparative evaluations demonstrate that SCoT significantly outperforms existing baselines in terms of resilience and response accuracy. Specifically, the results reveal that SCoT achieves a near-zero attack success rate, effectively generalizing to unseen attacks such as those attempting to suppress refusals (wei2024jailbroken). Moreover, it maintains robust general capabilities across standard operational scenarios, underscoring its superior adaptation to evolving security challenges and highlighting its sustained effectiveness.
To summarize, our contributions are three-fold:
  • A novel defense methodology. We propose a new defense approach, SCoT  which leverages reasoning-based analysis to defend against the most sophisticated and powerful attacks against LLMs, fundamentally diverging from traditional defenses.
  • Strong defense performance. Through comprehensive experiments, we show that SCoT significantly outperforms state-of-the-art defense strategies, offering superior robustness against out-of-distribution and adversarial attacks.
  • Model and resources release. We open-source our trained model and accompanying resources to facilitate further research and encourage collaboration within the AI safety community.
2 Background and Preliminary
@@ 2.1 Large Language Models
A large language model $\mathcal{M}$ generates human-like text by predicting words iteratively: $\hat{x_{i}}=\mathcal{M}(x_{i}\mid x_{1},\dots,x_{i-1})$ given the preceding sequence $(x_{1},\dots,x_{i-1})$. State-of-the-art LLMs leverage transformer architectures (vaswani2017attention) and are trained on large-scale corpora (openai2024gpt4technicalreport). Despite their capabilities, they remain vulnerable to jailbreak attacks. This work focuses on strengthening LLMs’ resilience against such attacks. Specifically, let $\mathcal{M}(q)$ denote the model’s response to a harmful query $q$. Effective safety mechanisms should ensure that the model consistently produces a refusal or a non-harmful response.
Figure 2: An overview of the Safety-Chain-of-Thought Methodology.
@@ 2.2 Jailbreak Attacks
Linguistic Manipulation. The key idea of this strategy is to alter the linguistic tone of the harmful question $q$ to evade LLM’s safety checking. Examples of linguistic manipulation include translating text into low-resource languages (deng2024multilingualjailbreakchallengeslarge), employing slang (xie2024sorrybenchsystematicallyevaluatinglarge), performing ASCII transformations, using Base64 encoding (wei2024jailbroken), and introducing intentional misspellings. These techniques can effectively bypass safety mechanisms by transforming input tokens into scenarios that appear out-of-distribution.
Contextual Manipulation. This strategy alters $q$ to $q^{\prime}$ by incorporating specific contextual elements like background information or persuasive language. Examples of contextual manipulation include adding role-play scenarios, evidence-based persuasion, or logical appeal that are designed to manipulate model behavior (xie2024sorrybenchsystematicallyevaluatinglarge; wei2024jailbroken). Such attacks typically involve meticulously crafted, human-written prompts that strategically influence the model’s responses. They exploit the model’s vulnerabilities by either prompting it to have a competing objective to ignore system instructions or by presenting inputs that the model does not recognize as harmful, thereby bypassing the established safety mechanisms (wei2024jailbroken).
Adaptive Attack. An adaptive attack views the challenge as an optimization problem, iteratively modifying $q$ to $\{q_{1},q_{2}\dots,q_{n}\}$ via optimizing a fitness function that indicates the likelihood of an affirmative answer. GCG (zou2023universal) defines the fitness function as the loss of affirmative responses with respect to the input, utilizing fuzzing techniques to refine $q_{i}$ until a successful output is achieved or the query budget is exhausted. AutoDAN (liu2024autodan) uses evolutionary algorithms to generate mutations for input modification. PAIR (chao2024jailbreakingblackboxlarge) diverges by using an external LLM to propose modifications based on the attack history $Q=\{q_{1},q_{2},\dots,q_{i}\}$. Those attack take advantage of both the competing objective and out-of-distribution simultaneously.
2.3 Defense
Existing defense strategies against jailbreak attacks on LLMs include refusal training, such as Llama3-8b-Instruct (schulman2017proximalpolicyoptimizationalgorithms; bai2022traininghelpfulharmlessassistant), adversarial training (AdvTrain) (mazeika2024harmbenchstandardizedevaluationframework), and Circuitbreaker (zou2024circuitbreaker). Refusal training, particularly when combined with techniques like RLHF and PPO (schulman2017proximalpolicyoptimizationalgorithms), empowers LLMs to reject unsafe prompts by reinforcing ethical decision-making during training. AdvTrain, motivated by adversarial training in computer vision, enhances models by fine-tuning with adversarial examples to help them recognize and resist harmful manipulations. On the other hand, Circuitbreaker aims to prevent the generation of harmful content by actively removing dangerous knowledge from the model during processing.
@@ 3 Our Approach
To address the aforementioned challenges, we propose a novel jailbreak defense method, Safety Chain-of-Thought (SCoT), summarized in Figure 2. Unlike traditional refusal training techniques that immediately block responses upon detecting harmful content (bai2022traininghelpfulharmlessassistant), SCoT requires the model to proactively analyze the harmful intent behind user requests before generating responses. This approach contains three key stages. First, we ❶ enhance the complexity and diversity of adversarial scenarios through question evolution (Section 3.1), which expands harmful questions through jailbreak mutations. We then ❷ establish a structured cognitive process for analyzing requests through malicious intent abstraction and safety regulation assessment for both harmful and benign question dataset (Section 3.2). Finally, we apply ❸ supervised fine-tuning on the newly developed safety reasoning dataset to enhance model’s broader reasoning capabilities while reinforcing its ability to resist jailbreak attempts (Section 3.3).
@@ 3.1 Question Evolution
As described in Section 2.2, jailbreak attacks typically employ two main strategies: (1) linguistic manipulation, which subtly alters the question’s phrasing with different tones, and (2) contextual manipulation, which includes distracting content to provoke affirmative or unintended responses. Inspired by the advancements in complex training scenarios demonstrated by WizardLM (xu2024wizardlm; luo2024wizardcoder), we aim to enhance the model’s reasoning capabilities by introducing more intricate questions. We develop complexity through two specific methods: depth evolution, which focuses on linguistic nuances, and breadth evolution, which incorporates distracting elements. This strategy is designed to counteract the effectiveness of adversarial manipulations by improving the model’s ability to discern and respond to nuanced intents within adversarial inputs.
Linguistic Manipulation. Linguistic manipulation involves subtly altering the phrasing of questions to bypass model defenses while retaining their original intent. In response, we enrich the question $q$ with diverse linguistic styles, specifically slang and uncommon dialects, to enhance the attack robustness against detection.
We implement this by utilizing In-Context Learning with a contemporary LLM, supported by a carefully curated set of demonstrations. To ensure the generation of diverse stylistic variants, we need to select demonstrations with different styles. We employ SorryBench (xie2024sorrybenchsystematicallyevaluatinglarge), which includes a rich corpus of slang and uncommon dialect texts. We begin by extracting semantic embeddings from the dataset entries using SentenceBERT (reimers2019sentencebertsentenceembeddingsusing). We then apply k-means clustering to these embeddings with $k$ clusters and select the centers of these clusters as our demonstrations. To further enhance the style diversity in the model’s outputs, we design explicit prompts for the LLM that encourage the generation of responses in varied styles, rather than strictly replicating the demonstrated examples. Details of the full prompt structure and the selected demonstrations are provided in Appendix A.2 and Appendix A.3. An example of our question evolution with linguistic manipulation is in Figure 3.
@@
We adopt a methodology akin to that for linguistic manipulations, where we carefully select a set of representative and diverse examples from the SorryBench (xie2024sorrybenchsystematicallyevaluatinglarge) subset with different persuasion techniques. These examples serve as demonstrations for In-Context Learning with our LLM, enabling to generate questions that are intricately complex and less predictable. The detailed prompt is in Appendix A.2. Due to the page limit, an example of our question evolution for contextual manipulation is in Appendix A.4.
@@ 3.2 Cognitive Reasoning Construction
Given a refusal training dataset $\mathcal{D}=\{(q,r)\}^{n}$ containing $n$ pairs of questions $q$ and refusal responses $r$, our goal is to enhance the refusal responses by incorporating safety reasoning in the form of a Chain-of-Thought guided by safety regulations. Leveraging this classification, we craft prompts for leading LLMs, such as GPT4, to generate responses in a structured three-stage format: (1) analyzing the intent of user requests, (2) explaining why the request does or does not align with one or more identified harmful categories, and (3) issuing a refusal response. Details of these prompts are given in Appendix A.3. The safety CoT is shown as follows:
#### Verify Request
I need to verify the dangerousness of the
users request to [summary] with following
categories before generating the response.
* harassment/threatening: [reason]
* hate/threatening: [reason]
* illicit/violent: [reason]
* self-harm/intent/instructions: [reason]
* sexual/minors: [reason]
* violence/graphic: [reason]
* misinformation: [reason]
* privacy/sensitive: [reason]
Therefore, the users request likely
involves a clear reference to [categories],
which could lead to unsafe behavior.
Therefore, I need to reject the request.
#### Response
I am sorry, but I cannot help with that.
Here, [summary] summarizes each question, encouraging the LLM to analyze the query intent first, especially when confronted with distracting content. The placeholder [reason] is replaced with “Yes” along with a valid explanation if applicable, or “No” otherwise. Finally, [categories] list the relevant categories identified. In the reasoning construction process, we primarily identify eight categories of potentially harmful content, i.e., harassment, hate speech, illicit/violent content, self-harm, sexual content, violence, misinformation, and privacy violations, by following OpenAI’s safety guidelines  and Google’s regulation framework .

### v2

# Enhancing Model Defense Against Jailbreaks with Proactive Safety Reasoning

@@ Abstract
Large language models (LLMs) are vital for a wide range of applications yet remain susceptible to jailbreak threats, which could lead to the generation of inappropriate responses. Conventional defenses, such as refusal and adversarial training, often fail to cover corner cases or rare domains, leaving LLMs still vulnerable to more sophisticated attacks. We propose a novel defense strategy, Safety Chain-of-Thought (SCoT), which harnesses the enhanced reasoning capabilities of LLMs for proactive assessment of harmful inputs, rather than simply blocking them. SCoT augments any refusal training datasets to critically analyze the intent behind each request before generating answers. By employing proactive reasoning, SCoT enhances the generalization of LLMs across varied harmful queries and scenarios not covered in the safety alignment corpus. Additionally, it generates detailed refusals specifying the rules violated. Comparative evaluations show that SCoT surpasses existing defenses, reducing vulnerability to out-of-distribution issues and adaptive attacks while maintaining strong general capabilities. Our implementation is available at https://github.com/xianglinyang/SafetyReasoningDataEvol.
@@ 1 Introduction
Large language models (LLMs) exhibit exceptional capabilities, enabling their wide applications in fields such as education [56], programming [34], and everyday tasks. However, their powerful nature also introduces risks, as they can inadvertently generate harmful instructions or inappropriate content, leading to unsafe or illegal outcomes. For instance, prior research has shown that malicious users can compromise LLMs through adversarial techniques such as jailbreaks [64, 27, 40], which bypass safety restrictions and prompt models to generate harmful outputs. These outputs may include instructions for breaching computer systems [64] or facilitating unauthorized access to copyrighted materials [30]. Such vulnerabilities not only pose ethical concerns but can also result in substantial real-world consequences, including financial loss, privacy violations, and legal liabilities for both users and organizations deploying these systems. Consequently, ensuring robust safety mechanisms for LLMs has become a critical priority for responsible AI development and deployment. Addressing safety challenges is not only critical to preventing potential harm but also essential to fostering trust and acceptance of these technologies in society.
A common way to achieve safety defense is safety alignment or refusal training [5], where LLMs are trained to reject harmful instructions. However, these strategies fail to defend against increasingly sophisticated attacks as their design and training recipes are still vulnerable to these emerging threats. These attacks often take advantage of two main weaknesses of the refusal training: (1) out-of-distribution scenarios, where the harmful input is not covered by safety training datasets but is understandable for the LLM due to exposure during the pretraining stage, and (2) competing objectives, where attackers could add harmful suffixes or introduce distracting role-play content to harmful questions to manipulate the model into starting its response positively [45, 3]. These methods exploit the way LLMs are trained and predict text, leading to inappropriate outputs. Even advanced models like GPT-4 [32] and Claude 3.5 Sonnet [4] remain vulnerable to these sophisticated methods, underscoring the urgent need for more adaptive and robust defense mechanisms.
We envision that human’s thinking process of responding to a given harmful query is to assess whether it is appropriate to give the answer in mind and then provide the response to the answer or refusal. Motivated by this cognitive process humans approach complex decisions and the recent advancements in the reasoning abilities of LLMs, we propose a novel defense strategy that fundamentally diverges from traditional approaches. This method, which we name SCoT: Safety Chain-of-Thought, instructs the model to first analyze the potential harmfulness of the request before giving the corresponding response. As depicted in Figure 1, this reasoning-based approach allows the model to engage in a “thinking process” to categorize the request into predefined violations, enabling it to generalize effectively to different variants of harmful questions and out-of-distribution scenarios that are not explicitly covered in the safety training corpus. Moreover, this proactive reflective mechanism naturally defends against attacks designed to elicit affirmative outputs.
@@
By requiring the model to evaluate the request’s intent before responding, our SCoT mitigates the risk of manipulation, ensuring a higher degree of robustness and adaptability compared to traditional defenses.
We rigorously evaluate SCoT across a comprehensive range of attack scenarios, encompassing 26 attack types. SCoT is benchmarked against safety-aligned models and some of the most advanced state-of-the-art defense strategies as outlined in recent studies [63, 30]. Comparative evaluations demonstrate that SCoT outperforms existing baselines in terms of harmful request resilience and general capabilities maintenance. Specifically, the results reveal that SCoT achieves a near-zero attack success rate, effectively generalizing to unseen attacks such as those attempting to suppress refusals [45]. Moreover, it maintains robust general capabilities across standard operational scenarios, underscoring its superior adaptation to evolving security challenges and highlighting its sustained effectiveness.
To summarize, our contributions are: (1) A novel defense methodology. We propose a new defense approach, SCoT  which leverages reasoning-based analysis to defend against the most sophisticated and powerful attacks against LLMs, fundamentally diverging from traditional defenses. (2) Strong defense performance. Through comprehensive experiments, we show that SCoT outperforms state-of-the-art defense strategies, offering superior robustness against out-of-distribution and adversarial attacks. (3) Model and resources release. We open-source our trained model and accompanying resources to facilitate further research and encourage collaboration within the AI safety community.
2 Background and Preliminaries
@@ 2.1 Large Language Models
A large language model $\mathcal{M}$ generates human-like text by predicting words iteratively: $\hat{x_{i}}=\mathcal{M}(x_{i}\mid x_{1},\dots,x_{i-1})$ given the preceding sequence $(x_{1},\dots,x_{i-1})$. State-of-the-art LLMs leverage transformer architectures [43] and are trained on large-scale corpora [32]. Despite their capabilities, they remain vulnerable to jailbreak attacks. This work focuses on strengthening LLMs’ resilience against attacks. Specifically, let $\mathcal{M}(q)$ denote the model’s response to a harmful query $q$. Effective safety mechanisms should ensure that the model consistently produces a refusal or a non-harmful response.
@@ 2.2 Jailbreak Attacks
Linguistic Manipulation. The key idea of this strategy is to alter the linguistic tone of the harmful question $q$ to evade LLM’s safety checking. Examples of linguistic manipulation include translating text into low-resource languages [11], employing slang [48], performing ASCII transformations, using Base64 encoding [45], and introducing intentional misspellings. These techniques can effectively bypass safety mechanisms by transforming input tokens into scenarios that appear out-of-distribution.
Contextual Manipulation. This strategy alters $q$ to $q^{\prime}$ by incorporating specific contextual elements like background information or persuasive language. Examples of contextual manipulation include adding role-play scenarios, evidence-based persuasion, technical terms [13] or logical appeal that are designed to manipulate model behavior [48, 45]. Such attacks typically involve meticulously crafted, human-written prompts that strategically influence the model’s responses. They exploit the model’s vulnerabilities by either prompting it to have a competing objective to ignore system instructions or by presenting inputs that the model does not recognize as harmful, thereby bypassing the established safety mechanisms [45].
Adaptive Attack. An adaptive attack views the jailbreak challenge as an optimization problem, iteratively refining $q$ into a sequence $\{q_{1},q_{2}\dots,q_{n}\}$, guided by a fitness function that estimates the likelihood of each $q_{i}$ eliciting an affirmative response from the target model [12, 23, 8, 54, 31]. GCG [64] defines the fitness function as the loss of affirmative responses with respect to the input, utilizing fuzzing techniques to refine $q_{i}$ until a successful output is achieved or the query budget is exhausted. AutoDAN [27] uses evolutionary algorithms to generate mutations for input modification. PAIR [6] diverges by using an external LLM to propose modifications based on the attack history $Q=\{q_{1},q_{2},\dots,q_{i}\}$. Those attacks take advantage of both the competing objective and out-of-distribution simultaneously.
2.3 Defense Strategies
Existing defense strategies against jailbreak attacks on LLMs include refusal training, such as Llama3-8b-Instruct [39, 5], adversarial training (R2D2[30], and RR [63]. Refusal training, particularly when combined with techniques like RLHF and PPO [39], empowers LLMs to reject unsafe prompts by reinforcing ethical decision-making during training. R2D2, motivated by adversarial training in computer vision, enhances models by fine-tuning with adversarial examples to help them recognize and resist harmful manipulations. On the other hand, RRaims to prevent the generation of harmful content by actively removing dangerous knowledge from the model during processing.
@@ 3 Our Approach
To address the aforementioned challenges, we propose a novel jailbreak defense method, Safety Chain-of-Thought (SCoT), summarized in Figure 2. Unlike traditional refusal training techniques that immediately block responses upon detecting harmful content [5], SCoT requires the model to proactively analyze the harmful intent behind user requests before generating responses. This approach contains three key stages. First, we ❶ enhance the complexity and diversity of adversarial scenarios through question evolution (Section 3.1), which expands harmful questions through jailbreak mutations. We then ❷ establish a structured cognitive process for analyzing requests through malicious intent abstraction and safety regulation assessment for both harmful and benign question dataset (Section 3.2). Finally, we apply ❸ supervised fine-tuning on the newly developed safety reasoning dataset to enhance model’s broader reasoning capabilities while reinforcing its ability to resist jailbreak attempts (Section 3.3).
Figure 2: An overview of the Safety-Chain-of-Thought Methodology.
@@ 3.1 Question Evolution
As described in Section 2.2, jailbreak attacks typically employ two main strategies: (1) linguistic manipulation, which subtly alters the question’s phrasing with different tones, and (2) contextual manipulation, which includes distracting content to provoke affirmative or unintended responses. Inspired by the advancements in complex training scenarios demonstrated by WizardLM [50, 29], we aim to enhance the model’s reasoning capabilities by introducing more intricate questions. We develop complexity through two specific methods: depth evolution, which focuses on linguistic nuances, and breadth evolution, which incorporates distracting elements. This strategy is designed to counteract the effectiveness of adversarial manipulations by improving the model’s ability to discern and respond to nuanced intents within adversarial inputs.
Linguistic Manipulation. Linguistic manipulation involves subtly altering the phrasing of questions to bypass model defenses while retaining their original intent. In response, we enrich the question $q$ with diverse linguistic styles, specifically slang and uncommon dialects, to enhance the attack robustness against detection. An example of our question evolution with linguistic manipulation is in Figure 3.
@@
We implement linguistic mutation using a powerful LLM such as GPT-4, guided by In-Context Learning (ICL) with a carefully curated set of demonstrations. These demonstrations are drawn from SorryBench [48], a rich corpus featuring slang and uncommon dialects. To ensure the generation of diverse stylistic variants, we need to select demonstrations with different styles. Specifically, we first extract semantic embeddings from the SorryBench entries using SentenceBERT [37], then apply k-means clustering to group them into $k$ clusters. We use the cluster centers as our final set of demonstrations. To further encourage stylistic variation, our prompts are designed to elicit diverse output styles rather than direct imitation of the demonstrations. Details of the demonstration selection algorithm are provided in Appendix A.1.
@@
We adopt a methodology akin to that for linguistic manipulations, where we carefully select a set of representative and diverse examples from the SorryBench [48] subset with different persuasion techniques. These examples serve as demonstrations for ICL with our LLM, enabling to generate questions that are intricately complex and less predictable. The detailed prompt is in Appendix B.2. An example of our question evolution for contextual manipulation is in Appendix A.2.
@@ 3.2 Cognitive Reasoning Construction
Given a refusal training dataset $\mathcal{D}=\{(q,r)\}^{n}$ containing $n$ pairs of questions $q$ and refusal responses $r$, our goal is to enhance the refusal responses by incorporating safety reasoning in the form of a Chain-of-Thought guided by safety regulations. As illustrated in Figure 4, each response is structured into three stages: (1) analyzing the intent behind the user request, (2) explaining whether and why the request aligns with one or more harmful content categories, and (3) issuing a refusal statement.
Figure 4: Safe CoT template with a structured three-stage format.
The [summary] placeholder captures the user’s intent, helping the LLM focus on the core request even when confronted with distracting or misleading phrasing. The [reason] placeholder is replaced with "Yes" and a justification if the request falls under harmful content, or "No" otherwise. The [harmful categories] field lists the specific safety categories identified during the analysis. We primarily consider eight categories of potentially harmful content—harassment, hate speech, illicit or violent activity, self-harm, sexual content, violence, misinformation, and privacy violations—based on OpenAI’s safety guidelines [33].
In practice, we use the leading LLMs such as GPT-4 to generate the intent summary and identify harmful categories for each request. We then filter out inaccurate or none outputs through post-processing. Details of the prompting strategy are provided in Appendix B.2.